# The computational power of discrete chemical reaction networks with bounded executions

### v2

@@ 6.4 Feedforward CRNs
Reaction-feedforward CRNs are significant in the sense that many continuous real-valued CRNs computing numerical-valued functions (where the count of some species $Y$ is interpreted as the output, e.g., $2X\to Y$ computes $f(x)=\lfloor x/2\rfloor$) can be computed by reaction-feedforward CRNs [chen2023rate].666 The definition of feedforward in reference [chen2023rate] is different from the definition given here, being based on an ordering of species rather than reactions. However, it is straightforward to verify by inspection that the CRNs given for the positive results of [chen2023rate] are reaction-feedforward according to Definition 6.26. Compared to general CRNs, reaction-feedforward CRNs are easy to analyze and prove correctness. One reason is that, if a reaction-feedforward CRN can reach terminal configuration from $\mathbf{x}$ at all, then it is execution bounded from $\mathbf{x}$.
There is a similar definition, called simply feedforward in [chen2023rate], based on ordering of species rather than reactions. We use the term species-feedforward to avoid confusion with Definition 6.26. We say a reaction $(\mathbf{r},\mathbf{p})$ produces a species $S$ if $\mathbf{p}(S)>\mathbf{r}(S)$, and it consumes $S$ if $\mathbf{r}(S)>\mathbf{p}(S)$.
@@
Although the term “linear potential function” was not used in [chen2023rate], it is shown in [chen2023rate, Lemma 4.8] that species-feedforward CRNs have a linear potential function (assigning weight $\frac{1}{K^{i}}$ to species $S_{i}$ for a suitably large constant $K$), thus are entirely execution bounded. The same is not always true of reaction-feedforward CRNs, for example $X\mathop{\rightarrow}\limits 2X$ is reaction-feedforward but not execution bounded. However, we can use similar techniques to proofs used for so-called noncompetitive CRNs in [vasic2022programming] to show “reasonable” reaction-feedforward CRNs are execution bounded.
@@
In a reaction-feedforward CRN $\mathcal{C}$, if there is a terminal configuration $\mathbf{c}_{\mathbf{i}}$ reachable from initial configuration $\mathbf{i}$, then $\mathbf{c}_{\mathbf{i}}$ is reached by every sufficiently long execution from $\mathbf{i}$. Furthermore, all of these executions are permutations of the same number of each reaction type. In particular, $\mathcal{C}$ is execution bounded from $\mathbf{i}$.
@@
Let $P$ be the execution leading from $\mathbf{i}$ to $\mathbf{c}_{\mathbf{i}}$. Consider any execution $Q$ with $|Q|>|P|$. By the pigeonhole principle, $Q$ must involve more occurrences of some reaction $r$ than $P$ does. By Lemma 6.28, this would imply that $\mathbf{c}_{\mathbf{i}}$ is not terminal, which contradicts the premise that $\mathbf{c}_{\mathbf{i}}$ is terminal. Therefore, no execution $Q$ can be longer than $P$. Consider any execution $Q$ where $|Q|=|P|$. $Q$ must be a permutation of $P$, as any deviation resulting in more of any reaction would, by the pigeonhole principle, lead to a contradiction of the terminality of $\mathbf{c}_{\mathbf{i}}$. To address the possibility of a shorter terminal execution, consider any execution $Q$ with $|Q|<|P|$. There must be some reaction $r$ occurring more frequently in $P$ than in $Q$, and by Lemma 6.28, $Q$ cannot reach a terminal configuration.
As noted, in the model of continuous CRNs, it is known that all the functions that can be stably computed (the continuous, piecewise linear functions) can be stably computed by reaction-feedforward CRNs [chen2023rate]. In contrast, with discrete CRNs computing predicates, we show that reaction-feedforward CRNs cannot stably decide all semilinear sets by giving two counterexamples, showing that reaction-feedforward CRDs can decide neither “most” mod sets (6.31) nor “most” threshold sets (6.32). Specifically, we chose the parity and majority predicate as our counterexamples, although the techniques generalize to more complex mod and threshold sets, e.g., $[X_{1}+2X_{2}\equiv 3\mod 5?]$.
@@ 7 Conclusion
A key question remains open: Can execution bounded CRNs compute semilinear functions and predicates within polylogarithmic time? Angluin, Aspnes and Eisenstat introduced a fast population protocol that simulates a register machine with high probability. This protocol can perform standard operations like comparison, addition, subtraction, and multiplication and division by constants in $O(\log^{5}n)$ time [AngluinAE2008Fast]. Chen, Doty and Soloveichik applied this construction to CRNs in [Chen2012DeterministicFunction], showing that semilinear functions can be computed by CRNs without error in expected polylogarithmic time in the kinetic model. Central to their success in both cases is the “phase clock”, which generates a clock signal to indicate the probable completion of an epidemic style chain reaction and orders more recent instructions to overwrite older ones. This clock is inherently unbounded in its execution, cycling through $m$ stages.

### v3

@@ 6.4 Feedforward CRNs
Reaction-feedforward CRNs are significant in the sense that many continuous real-valued CRNs computing numerical-valued functions (where the count of some species $Y$ is interpreted as the output, e.g., $2X\to Y$ computes $f(x)=\lfloor x/2\rfloor$) can be computed by reaction-feedforward CRNs [4].666 The definition of feedforward in reference [4] is different from the definition given here, being based on an ordering of species rather than reactions. However, it is straightforward to verify by inspection that the CRNs given for the positive results of [4] are reaction-feedforward according to Definition 6.26. Compared to general CRNs, reaction-feedforward CRNs are easy to analyze and prove correctness. One reason is that, if a reaction-feedforward CRN can reach terminal configuration from $\mathbf{x}$ at all, then it is execution bounded from $\mathbf{x}$.
There is a similar definition, called simply feedforward in [4], based on ordering of species rather than reactions. We use the term species-feedforward to avoid confusion with Definition 6.26. We say a reaction $(\mathbf{r},\mathbf{p})$ produces a species $S$ if $\mathbf{p}(S)>\mathbf{r}(S)$, and it consumes $S$ if $\mathbf{r}(S)>\mathbf{p}(S)$.
@@
Although the term “linear potential function” was not used in [4], it is shown in [4, Lemma 4.8] that species-feedforward CRNs have a linear potential function (assigning weight $\frac{1}{K^{i}}$ to species $S_{i}$ for a suitably large constant $K$), thus are entirely execution bounded. The same is not always true of reaction-feedforward CRNs, for example $X\mathop{\rightarrow}\limits 2X$ is reaction-feedforward but not execution bounded. However, we can use similar techniques to proofs used for so-called noncompetitive CRNs in [18] to show “reasonable” reaction-feedforward CRNs are execution bounded.
@@
In a reaction-feedforward CRN $\mathcal{C}$, if there is a terminal configuration $\mathbf{c}_{\mathbf{i}}$ reachable from initial configuration $\mathbf{i}$, then $\mathbf{c}_{\mathbf{i}}$ is reached by every sufficiently long execution from $\mathbf{i}$, so $\mathbf{c}_{\mathbf{i}}$ is the only terminal configuration reachable from $\mathbf{i}$. Furthermore, all of these executions are permutations of the same number of each reaction type. In particular, $\mathcal{C}$ is execution bounded from $\mathbf{i}$.
@@
Let $P$ be the execution leading from $\mathbf{i}$ to $\mathbf{c}_{\mathbf{i}}$. Let $Q$ be any execution from $\mathbf{i}$.
We first claim that $|Q|\leq|P|$; suppose otherwise the sake of contradiction. By the pigeonhole principle, $Q$ must have more of some reaction than $P$ does. By Lemma 6.28, this would imply that $\mathbf{c}_{\mathbf{i}}$ is not terminal, a contradiction. Therefore, no execution $Q$ can be longer than $P$; in particular $\mathcal{C}$ is execution bounded from $\mathbf{i}$.
If $|Q|=|P|$, we claim $Q$ must be a permutation of $P$. Supposing otherwise, by the pigeonhole principle, $Q$ would have more of some reaction than $P$, again with Lemma 6.28 implying that $\mathbf{c}_{\mathbf{i}}$ is not terminal, a contradiction.
Finally, suppose $|Q|<|P|$. By pigeonhole some reaction occurs more frequently in $P$ than in $Q$, so by Lemma 6.28, $Q$ cannot reach a terminal configuration. Therefore all executions from $\mathbf{i}$ to $\mathbf{c}_{\mathbf{i}}$ are permutations of $P$.
As noted, in the model of continuous CRNs, it is known that all the functions that can be stably computed (the continuous, piecewise linear functions) can be stably computed by reaction-feedforward CRNs [4]. In contrast, with discrete CRNs computing predicates, we show that reaction-feedforward CRNs cannot stably decide all semilinear sets by giving two counterexamples, showing that reaction-feedforward CRDs can decide neither “the simplest” mod set (6.31) nor “the simplest” threshold set (6.32). Specifically, we chose the parity and majority predicate as our counterexamples, although the techniques generalize to more complex mod and threshold sets, e.g., $[X_{1}+2X_{2}\equiv 3\mod 5?]$.
@@ 7 Conclusion
A key question remains open: Can execution bounded CRNs compute semilinear functions and predicates within polylogarithmic time? Angluin, Aspnes and Eisenstat [2] introduced a fast population protocol that simulates a register machine with high probability, and can be made probability 1 with semilinear predicates. However, this construction seems inherently unbounded in executions.